OpenStudy (sette):

solve for x in the proportion x 1 ---- = ---- x+12 x-1 answers are not exact. they are rounded. a: x=-6.02 and x=5.02 b: x=-3.46 and x=3.46 c: x=-3.14 and x=4.14 d: x=-2.61 and x=4.61

OpenStudy (anonymous):

solve this equation for x: \[x(x-1) - (x+12) = 0\] which reduces to a quadratic: \[x^2 - 2x - 12 = 0\] which we use the formula to solve: \[x=\frac{2 \pm \sqrt{52}}{2}\] \[x=\frac{2 \pm 2\sqrt{13}}{2}\] \[x=1 \pm \sqrt{13} = 1 \pm 3.6055\] \[ x = 4.61, x = -2.61\]

OpenStudy (anonymous):

so answer is d

OpenStudy (sette):

how did you get x(x-1)-(x+12) ?

OpenStudy (sette):

is it like a certain equation ? because to solve for proportions and ratios i usually cross multiply

OpenStudy (anonymous):

Take your starting equation: \[\frac{x}{x+12} = \frac{1}{x-1}\] multiply both sides by (x-1) to give \[\frac{x(x-1)}{x+12} = 1\] then multiply both sides by (x+12) \[x(x-1) = x+12\] so we have removed the fractions. Now just take (x+12) to the other side: \[x(x-1) - (x+12) = 0\]

OpenStudy (sette):

oh okay ! i see . thank you(: